OpenStudy (anonymous):

How are the sin 30° and the cos 60° related? Explain using complete sentences.

OpenStudy (anonymous):

sin 30=cos 60

OpenStudy (anonymous):

both 1/2

OpenStudy (anonymous):

u really should learn trignometry more....

OpenStudy (anonymous):

How do i explain it.. and im trying its just really hard for me

OpenStudy (anonymous):

@mathmale could u help him explain this

OpenStudy (anonymous):

im a girl! lol

OpenStudy (anonymous):

oh sorry..u shud really add a display picture...anyway nice to meet u

OpenStudy (anonymous):

its ok , im going to!

OpenStudy (anonymous):

and you should really study trigo more. this is the very basics of it., you'll be in trouble if this goes on and you don't learn the fundamentals of it

OpenStudy (anonymous):

If you will look at the Unit circle, which i recommend you see now, you will observe that cos at 60 and sin at 30 would have the same points, furthermore, if you look at the trigonometric table which i recommend you learn, you'd observe the same. And when you have a graph, you'll see that where sin is positive, the quadrant opposite is exactly where the cos is positive. Hope it helps.

OpenStudy (mathmale):

the sine and cosine are (believe it or not) called "co-functions." One property of these "co-functions" is that sin A = cos (90 degrees - A). Note that 30 and 60 add up to 90. We call 30 and 60 "complementary angles," by virtue of their adding up to 90. If two angles add to 180, they are "supplementary angles," in case you were dying to know more than you have to to answer this particular homework problem. :) Therefore, sin 30 = cos (90-30) = cos 60.